OpenStudy (anonymous):

Five consecutive even integers have a sum of -20. Quantity A = The greatest of the 5 integers. Quantity B= 0. Is it: A) Quantity A is greater B) Quantity B is greater C) They are equal D) Not enough info to determine.

OpenStudy (inkyvoyd):

Ok.

OpenStudy (inkyvoyd):

x (greatest of the five) x+(x-2)+(x-4)+(x-6)+(x-8)=-20 Solve for x

OpenStudy (inkyvoyd):

Or, you could do a faster way

OpenStudy (inkyvoyd):

@RaphaelFilgueiras , I do believe that is unecessarily nasty.

OpenStudy (anonymous):

how you can you say that x cannot be even?

OpenStudy (inkyvoyd):

No. x is even.

OpenStudy (inkyvoyd):

x is even because it is the largest of the sequence, and is defined to be even.

OpenStudy (inkyvoyd):

any consecutive even number has a difference of 2 from the previous number.

OpenStudy (anonymous):

i read odd,but you don't have to put 2K?

OpenStudy (inkyvoyd):

Five consecutive even integers have a sum of -20 I assumed that x was the largest number. Then, x must be even, because I assumed it was part of the sequ3ence.

OpenStudy (anonymous):

i dont know is think yoy have to write this way (2k)+(2k+2)+(2K+4)+(2k+6)+(2k+8)=−20

OpenStudy (anonymous):

ok i got it!

OpenStudy (anonymous):

but if we have 4...

OpenStudy (anonymous):

@RaphaelFilgueiras, usually, in a proof involving even integers, you have to appeal to the fact that an even integer is of the form 2k for some integer k. In this case, since you START with a number that's defined to be even, it's enough to appeal to the fact that consecutive even integers from that number will have a difference of two. @inkyvoyd, you're method is good, but the problem defines the variable B= the highest integer, so I don't see any point in REdefining it. Just use: B +(B-2) +(B-4) +(B-6) +(B-8) = -20

OpenStudy (inkyvoyd):

I did. NOte my second post x (greatest of the five) x+(x-2)+(x-4)+(x-6)+(x-8)=-20 Solve for x

OpenStudy (inkyvoyd):

@SmoothMath

OpenStudy (anonymous):

Haha no. I'm saying you don't have to REDEFINE a variable that the problem already defines.

OpenStudy (anonymous):

It's just extra work and confuses the issue to have two variables named the same thing. B = highest even integer x = highest even integer We don't need both of these.